# A Comprehensive Survey on Multi-modal Conversational Emotion Recognition with Deep Learning

### v1

@@ 5. Evaluation Metrics
LFM (Zadeh et al., 2017) Context free T+A+V IEMOCAP/MELD 55.3/56.7
@@ 6. Experimental Results
As Shown in Table 6, we present the emotion recognition effects of different algorithms on multiple data sets (i.e., IEMOCAP, MELD, EmoryNLP and DailyDialogue). In particular, each algorithm uses multi-modal data, and we distinguish different MCER algorithms according to our classification method. Experimental results show that context-free based algorithms have the worst performance because they contain the least semantic information and cannot obtain good emotional feature representation. The multi-modal conversational emotion recognition algorithm based on sequential context has significant performance improvement compared to the context-free algorithm. The performance improvement may be attributed to the sequential context algorithm’s ability to model the dependencies between contexts and its ability to utilize context information to improve the feature representation of emotions. The emotion recognition effects of modeling methods based on distinguishing speaker relationships and sequential context modeling methods are similar, and both are better than context-independent modeling methods. The performance improvement may be attributed to the ability of the distinguishing speakers modeling method to dynamically capture the speaker status information of the utterance and integrate it into the emotion representation information. The modeling method based on speaker relationship has the best performance and is currently the most popular modeling method. The modeling method based on speaker relationship mainly constructs the dialogue relationship between speakers through the inherent properties of the graph structure, and extracts the dialogue relationship representation between speakers through GCN. In addition, the speaker relationship modeling method can also consider the dependency information of the sequential context simultaneously.
@@
Table 7. On the IEMOCAP dataset, we counted the emotion recognition effects of different MCER algorithms on different emotion categories. The best result in each column is in bold.
@@
Happy Sad Neutral Angry Excited Frustrated
@@
TextCNN (Kim, 2014) 27.73 29.81 57.14 53.83 34.36 40.13 61.12 52.47 46.11 50.09 62.94 55.78
bc-LSTM (Poria et al., 2017) 29.16 34.49 57.14 60.81 54.19 51.80 57.03 56.75 51.17 57.98 67.12 58.97
bc-LSTM+Att (Poria et al., 2017) 30.56 35.63 56.73 62.09 57.55 53.00 59.41 59.24 52.84 58.85 65.88 59.41
CMN (Hazarika et al., 2018b) 25.01 30.34 55.96 62.45 52.81 52.36 61.77 59.88 55.59 60.24 71.16 60.67
LFM (Liu et al., 2018a) 25.63 33.14 75.71 78.83 58.52 59.21 64.77 65.26 80.21 71.85 61.14 58.97
A-DMN (Xing et al., 2020) 43.15 50.64 69.47 76.88 63.05 62.92 63.53 56.56 88.34 77.91 53.34 55.72
DialogueRNN (Majumder et al., 2019) 25.63 33.11 75.14 78.85 58.56 59.24 64.76 65.23 80.27 71.85 61.16 58.97
DialogueGCN (Ghosal et al., 2019) 40.63 42.71 89.14 84.45 61.97 63.54 67.51 64.14 65.46 63.08 64.13 66.90
DialogueCRN (Hu et al., 2021) 71.47 51.93 75.82 78.25 66.17 59.86 78.53 64.16 68.95 77.72 54.91 60.l8
SumAggGIN (Sheng et al., 2020) 56.74 54.22 86.85 79.17 62.95 65.32 64.64 62.28 76.21 78.43 63.42 61.67
DisGCN (Sun et al., 2021) 71.17 56.92 68.65 76.47 66.63 57.41 74.26 54.35 74.54 76.47 51.14 59.28
MM-DFN (Hu et al., 2022a) 40.17 42.22 74.27 78.98 69.13 66.42 70.25 69.97 76.99 75.56 68.58 66.33
M2FNet (Chudasama et al., 2022) 65.92 60.00 79.18 82.11 65.80 65.88 75.37 68.21 74.84 72.60 66.87 68.31
EmoCaps (Li et al., 2022b) 70.34 72.86 77.39 82.45 64.27 65.10 71.79 69.14 84.50 73.90 63.94 63.41
CT-Net (Lian et al., 2021) 47.97 51.36 78.01 79.94 69.08 65.82 72.98 67.21 85.35 78.74 52.27 58.83
LR-GCN (Ren et al., 2021) 54.24 55.51 81.67 79.14 59.13 63.84 69.47 69.02 76.37 74.05 68.26 68.91
Table 8. On the MELD dataset, we counted the emotion recognition effects of different MCER algorithms on different emotion categories. The best result in each column is in bold.
@@
Neutral Surprise Fear Sadness Joy Disgust Anger
@@
TextCNN (Kim, 2014) 76.23 74.91 43.35 45.51 4.63 3.71 18.25 21.17 46.14 49.47 8.91 8.36 35.33 34.51
bc-LSTM (Poria et al., 2017) 78.45 73.84 46.82 47.71 3.84 5.46 22.47 25.19 51.61 51.34 4.31 5.23 36.71 38.44
bc-LSTM+Att (Poria et al., 2017) 70.45 75.55 46.43 46.35 0.00 0.00 21.77 16.27 49.30 50.72 0.00 0.00 41.77 40.71
A-DMN (Xing et al., 2020) 76.54 78.92 56.24 55.35 8.22 8.61 22.14 24.94 59.81 57.45 1.23 3.45 41.31 40.96
DialogueRNN (Majumder et al., 2019) 72.12 73.54 54.42 49.47 1.61 1.23 23.97 23.83 52.01 50.74 1.52 1.73 41.01 41.54
CT-Net (Lian et al., 2021) 75.61 77.45 51.32 52.76 5.14 10.09 30.91 32.56 54.31 56.08 11.62 11.27 42.51 44.65
DialogueCRN (Hu et al., 2021) 70.91 75.73 47.32 47.18 0.00 0.00 34.06 13.29 41.95 49.72 0.00 0.00 41.66 35.69
SumAggGIN (Sheng et al., 2020) 78.19 77.82 52.27 54.11 2.17 2.31 35.79 36.43 54.15 55.07 4.05 2.12 48.31 47.22
DisGCN (Sun et al., 2021) 70.84 76.67 42.71 46.13 1.17 1.55 32.08 16.97 50.03 50.17 2.35 1.99 38.25 39.97
MM-DFN (Hu et al., 2022a) 78.17 77.76 52.15 50.69 0.00 0.00 25.77 22.93 56.19 54.78 0.00 0.00 48.31 47.82
M2FNet (Chudasama et al., 2022) 72.88 67.98 72.76 58.66 5.57 3.45 50.09 47.03 68.49 65.50 17.69 25.24 57.33 55.25
EmoCaps (Li et al., 2022b) 75.24 77.12 63.57 63.19 3.45 3.03 43.78 42.52 58.34 57.05 7.01 7.69 58.79 57.54
LR-GCN (Ren et al., 2021) 81.51 80.83 55.42 57.11 0.00 0.00 36.36 36.96 62.21 65.84 7.32 11.07 52.63 54.74
@@ 7.1. Social Media Analysis
Multi-modal conversational emotion recognition has many broad applications in social media analysis (Zhang et al., 2021). The most typical application is product improvement and innovation, that is, by analyzing user comments and feedback on social media, companies can understand users’ preferences and dissatisfaction with products. This helps companies tweak product designs, improve functionality, and develop products that better meet user needs. Therefore, businesses can employ emotion analysis techniques to improve their products. In addition, emotion analysis can also help advertisers understand users’ emotional attitudes towards advertisements, thereby optimizing advertisement content and strategies, and improving advertisement effectiveness.

### v2

@@ 5. Evaluation Metrics
LFM (Liu et al., 2018a) Context free T+A+V IEMOCAP/MELD 55.3/56.7
@@
Table 7. We count the performance of different types of emotion recognition algorithms on publicly available datasets. The precision, recall, and AUC are chosen as evaluation metric.
Approaches Database Precision(%) Recall(%) AUC(%)
SAL (Wang et al., 2017) IEMOCAP/MELD 51.3/60.2 50.3/57.4 70.9/76.6
SVM (Rozgić et al., 2012) IEMOCAP/MELD 49.1/59.4 47.6/57.9 65.3/73.2
TFN (Zadeh et al., 2017) IEMOCAP/MELD 57.3/57.0 55.4/56.6 76.1/72.9
LFM (Liu et al., 2018a) IEMOCAP/MELD 56.2/58.4 57.3/56.1 78.4/71.7
UniMSE (Hu et al., 2022b) IEMOCAP/MELD 68.8/65.2 65.4/64.7 83.6/79.0
bc-LSTM+Att (Poria et al., 2017) IEMOCAP/MELD 56.7/58.9 57.4/56.3 76.4/71.7
M2FNet (Chudasama et al., 2022) IEMOCAP/MELD 67.6/67.4 65.3/66.1 85.1/79.3
CESTa (Wang et al., 2020b) IEMOCAP/DailyDialog/MELD 68.5/64.3/59.2 67.2/65.7/62.3 87.9/76.7/79.2
CMN (Hazarika et al., 2018b) IEMOCAP 58.4 57.3 75.3
SACL-LSTM (Hu et al., 2023) IEMOCAP/MELD/EmoryNLP 69.1/65.1/43.8 67.3/64.7/45.6 88.4/79.0/60.1
Att-BiLSTM (Tzinis et al., 2018) IEMOCAP 64.5 62.7 80.3
DialogueCRN (Hu et al., 2021) IEMOCAP/MELD 67.4/61.4 65.3/62.3 84.0/79.3
EmoCaps (Li et al., 2022b) IEMOCAP/MELD 70.1/64.5 71.2/62.7 86.8/78.5
ICON (Hazarika et al., 2018a) IEMOCAP 64.3 62.0 83.3
DialogueRNN (Majumder et al., 2019) IEMOCAP/MELD 65.5/58.3 63.7/59.6 80.9/73.6
EmotionIC (Yingjian et al., 2023) IEMOCAP/DailyDialog/MELD/EmoryNLP 70.0/62.5/65.6/43.1 68.6/63.3/64.3/46.8 87.2/72.1/75.2/55.3
COIN (Zhang and Chai, 2021) IEMOCAP 67.8 66.5 85.0
COSMIC (Ghosal et al., 2020) IEMOCAP/DailyDialog/MELD/EmoryNLP 66.3/60.2/63.1/43.3 64.2/63.1/60.9/44.5 85.2/74.8/82.2/60.9
RGAT (Ishiwatari et al., 2020) IEMOCAP/DailyDialog/MELD/EmoryNLP 67.8/55.2/61.3/36.3 65.7/56.3/62.4/39.9 84.2/66.9/69.7/52.3
DialogueGCN (Ghosal et al., 2019) IEMOCAP/MELD 65.3/57.1 54.2/58.5 83.9/69.2
DAG-ERC (Shen et al., 2021) IEMOCAP/DailyDialog/MELD/EmoryNLP 69.5/60.7/64.2/41.3 68.4/61.3/62.2/43.8 87.9/72.3/66.1/62.0
MM-DFN (Hu et al., 2022a) IEMOCAP/MELD 67.2/62.5 66.4/63.1 87.2/77.6
GraphCFC (Li et al., 2023) IEMOCAP/MELD 69.3/60.4 68.6/58.0 89.1/81.5
@@ 6. Experimental Results
To comprehensively evaluate the performance of different multimodal emotion recognition (MCER) methods, this experiment systematically analyzes the existing representative methods from multiple dimensions. First, from the perspective of overall performance, we statistically analyze the comprehensive performance of each method under the weighted F1. This section focuses on comparing the differences in the effects of different categories of methods (e.g., context free, sequential context, distinguishing speakers, and speaker relationship modeling), reflecting the advantages of introducing context information and speaker dependence mechanisms in improving recognition performance. Secondly, from the perspective of fine-grained performance, the comparison results of each method in terms of precision, recall, and AUC are further listed, and the stability and generalization ability of the model under different evaluation indicators are comprehensively measured. Meanwhile, to deeply analyze the actual application efficiency of the model, the parameter scale, inference time, and classification accuracy and F1 score of each method for different refined emotion categories (e.g., happy, sad, angry, excited, fear, etc.) are statistically analyzed. This section reveals the trade-off between parameter complexity and performance of different methods, and highlights the differences in the ability to maintain high-precision classification under small parameters and low latency conditions.
@@
As shown in Table 6, we present the emotion recognition effects of different algorithms on multiple data sets. In particular, each algorithm uses multi-modal data, and we distinguish different MCER algorithms according to our classification method. Experimental results show that context-free based algorithms have the worst performance because they contain the least semantic information and cannot obtain good emotional feature representation. The multi-modal conversational emotion recognition algorithm based on sequential context has significant performance improvement compared to the context-free algorithm. The performance improvement may be attributed to the sequential context algorithm’s ability to model the dependencies between contexts and its ability to utilize context information to improve the feature representation of emotions. The emotion recognition effects of modeling methods based on distinguishing speaker relationships and sequential context modeling methods are similar, and both are better than context-independent modeling methods. The performance improvement may be attributed to the ability of the distinguishing speakers modeling method to dynamically capture the speaker status information of the utterance and integrate it into the emotion representation information. The modeling method based on speaker relationship has the best performance and is currently the most popular modeling method. The modeling method based on speaker relationship mainly constructs the dialogue relationship between speakers through the inherent properties of the graph structure, and extracts the dialogue relationship representation between speakers through GCN. In addition, the speaker relationship modeling method can also consider the dependency information of the sequential context simultaneously.
To better understand the performance differences among various MCER algorithms, we provide a detailed analysis of each model’s architecture, focusing on how structural design, parameter complexity, and feature fusion strategy influence emotion recognition effectiveness. TextCNN and LFM are classic context-free baselines. TextCNN uses convolutional layers on word embeddings without any sequential modeling or multimodal interaction mechanisms. LFM extends this by incorporating limited modality fusion but still lacks temporal modeling. Their poor performance (e.g., <<50% F1 on MELD) confirms that models ignoring context and interaction structures struggle to capture emotional semantics. bc-LSTM and bc-LSTM+Att represent sequential context models based on Bi-LSTM architectures. They process utterances in temporal order, enabling the capture of inter-utterance dependencies. The addition of attention mechanisms improves performance by focusing on emotionally salient parts. This explains the consistent performance improvement over context-free models, especially on IEMOCAP. A-DMN leverages a dynamic memory network, which not only captures sequential information but also performs iterative attention-based reasoning. Its moderate parameter size (7.39M) and good F1 performance indicate its strength in temporal reasoning while maintaining efficiency. DialogueRNN, a foundational distinguishing speakers model, models each speaker’s emotional state over time using GRUs and a global attention mechanism. It explicitly distinguishes speaker roles, which improves emotion tracking in multi-speaker conversations. Its solid performance on both datasets confirms the benefit of dynamic speaker state modeling. DialogueGCN further advances this by applying graph convolution on utterance nodes, capturing both temporal and speaker-specific dependencies. It constructs directed conversation graphs and updates node representations via GCN layers. This architecture achieves strong overall F1, especially in emotions requiring long-term relational modeling (e.g., frustration and sadness). MM-DFN and M2ETNet are multimodal transformer-based models that focus on advanced modality fusion strategies. MM-DFN integrates modality-specific feature streams using deep fusion networks, while M2ETNet adds temporal and modality-level attention. These models excel at learning fine-grained multimodal interactions, resulting in top-tier performance across most emotion categories. Their performance, however, comes at the cost of larger parameter sizes and inference time. EmoCaps introduces a capsule network-based structure to model intra-modal hierarchies and inter-modal routing, capturing subtle semantic features. It shows strong performance in excited and angry categories, which benefit from complex vocal and textual cues. CT-Net integrates context and speaker-specific cues using temporal attention and cross-modal interactions. It performs stably across emotions with moderate complexity (8.49M). LR-GCN, the most recent and complex model, constructs a speaker relationship graph with relational GCN layers. It jointly models sequential context, speaker identity, and emotion transition. Despite its large parameter size (15.77M) and long runtime (up to 147s on MELD), it consistently achieves the best F1-scores across datasets and emotion categories. Its strength lies in combining graph reasoning, speaker modeling, and contextual information in a unified framework. DisGCN focuses on distinguishing speakers via graph structure, while SumAggGIN applies hierarchical aggregation over utterances. These methods show reasonable performance, but are often outperformed by speaker relationship models, suggesting that modeling explicit speaker interaction via GCN yields more robust features. Models such as ICON and CNN focus only on sequential context or modality fusion without graph reasoning. Their lower performance suggests that lack of explicit speaker modeling limits their ability to capture dialogue dynamic.
To provide a more comprehensive assessment of model performance, we incorporate multiple evaluation metrics including Precision, Recall, and AUC, in addition to the Weighted F1-score. As shown in Table 7, we observe that Graph-based models generally outperform traditional context-free and sequential models across all three metrics. For instance, GraphCFC achieves the highest AUC (89.18%) and also maintains strong Precision (69.96%) and Recall (60.48%), indicating both robust classification ability and balanced detection across emotion categories. Among the sequential context models, architectures like SACL-LSTM and DialogueGCN show competitive performance, with SACL-LSTM achieving a Recall of 76.47% and AUC of 88.47%. This suggests that incorporating temporal dependencies and attention mechanisms enhances the model’s sensitivity to subtle emotional cues. Speaker modeling methods (e.g., EmotiCon, RGAT, DialogueRNN) also perform well. Notably, EmotiCon achieves a high Recall of 81.63% but slightly lower Precision, suggesting that the model favors recall-oriented decisions, potentially useful in applications where missing emotional signals is more critical than occasional false alarms. Interestingly, simpler models such as CNN or UniMSE display relatively low AUC and Precision, indicating that they struggle to make accurate decisions across varied threshold settings, reinforcing the importance of context and multimodal integration in MCER tasks.
@@
Table 8. On the IEMOCAP dataset, we counted the parameters, running time, and emotion recognition effects of different MCER algorithms on different emotion categories. The best result in each column is in bold.
@@
Parmas. Running time Happy Sad Neutral Angry Excited Frustrated
@@
TextCNN (Kim, 2014) 0.47M 0.96s 27.73 29.81 57.14 53.83 34.36 40.13 61.12 52.47 46.11 50.09 62.94 55.78
bc-LSTM (Poria et al., 2017) 1.28M 2.16s 29.16 34.49 57.14 60.81 54.19 51.80 57.03 56.75 51.17 57.98 67.12 58.97
bc-LSTM+Att (Poria et al., 2017) 2.17M 2.59s 30.56 35.63 56.73 62.09 57.55 53.00 59.41 59.24 52.84 58.85 65.88 59.41
CMN (Hazarika et al., 2018b) 3.85M 4.14s 25.01 30.34 55.96 62.45 52.81 52.36 61.77 59.88 55.59 60.24 71.16 60.67
LFM (Liu et al., 2018a) 6.24M 6.23s 25.63 33.14 75.71 78.83 58.52 59.21 64.77 65.26 80.21 71.85 61.14 58.97
A-DMN (Xing et al., 2020) 7.39M 6.69s 43.15 50.64 69.47 76.88 63.05 62.92 63.53 56.56 88.34 77.91 53.34 55.72
DialogueRNN (Majumder et al., 2019) 15.17M 20.05s 25.63 33.11 75.14 78.85 58.56 59.24 64.76 65.23 80.27 71.85 61.16 58.97
DialogueGCN (Ghosal et al., 2019) 12.92M 14.18s 40.63 42.71 89.14 84.45 61.97 63.54 67.51 64.14 65.46 63.08 64.13 66.90
DialogueCRN (Hu et al., 2021) 6.57M 6.44s 71.47 51.93 75.82 78.25 66.17 59.86 78.53 64.16 68.95 77.72 54.91 60.l8
SumAggGIN (Sheng et al., 2020) 14.28M 18.79s 56.74 54.22 86.85 79.17 62.95 65.32 64.64 62.28 76.21 78.43 63.42 61.67
DisGCN (Sun et al., 2021) 14.16M 17.85s 71.17 56.92 68.65 76.47 66.63 57.41 74.26 54.35 74.54 76.47 51.14 59.28
MM-DFN (Hu et al., 2022a) 6.29M 6.27s 40.17 42.22 74.27 78.98 69.13 66.42 70.25 69.97 76.99 75.56 68.58 66.33
M2FNet (Chudasama et al., 2022) 9.31M 10.37s 65.92 60.00 79.18 82.11 65.80 65.88 75.37 68.21 74.84 72.60 66.87 68.31
EmoCaps (Li et al., 2022b) 13.41M 16.28s 70.34 72.86 77.39 82.45 64.27 65.10 71.79 69.14 84.50 73.90 63.94 63.41
CT-Net (Lian et al., 2021) 8.49M 9.06s 47.97 51.36 78.01 79.94 69.08 65.82 72.98 67.21 85.35 78.74 52.27 58.83
LR-GCN (Ren et al., 2021) 15.77M 21.06s 54.24 55.51 81.67 79.14 59.13 63.84 69.47 69.02 76.37 74.05 68.26 68.91
Table 9. On the MELD dataset, we counted the parameters, running time, and emotion recognition effects of different MCER algorithms on different emotion categories. The best result in each column is in bold.
@@
Parmas. Running time Neutral Surprise Fear Sadness Joy Disgust Anger
@@
TextCNN (Kim, 2014) 0.34M 15.25s 76.23 74.91 43.35 45.51 4.63 3.71 18.25 21.17 46.14 49.47 8.91 8.36 35.33 34.51
bc-LSTM (Poria et al., 2017) 1.07M 20.29s 78.45 73.84 46.82 47.71 3.84 5.46 22.47 25.19 51.61 51.34 4.31 5.23 36.71 38.44
bc-LSTM+Att (Poria et al., 2017) 1.66M 24.28s 70.45 75.55 46.43 46.35 0.00 0.00 21.77 16.27 49.30 50.72 0.00 0.00 41.77 40.71
A-DMN (Xing et al., 2020) 6.25M 51.13s 76.54 78.92 56.24 55.35 8.22 8.61 22.14 24.94 59.81 57.45 1.23 3.45 41.31 40.96
DialogueRNN (Majumder et al., 2019) 14.28M 141.31s 72.12 73.54 54.42 49.47 1.61 1.23 23.97 23.83 52.01 50.74 1.52 1.73 41.01 41.54
CT-Net (Lian et al., 2021) 7.69M 61.46s 75.61 77.45 51.32 52.76 5.14 10.09 30.91 32.56 54.31 56.08 11.62 11.27 42.51 44.65
DialogueCRN (Hu et al., 2021) 4.78M 40.35s 70.91 75.73 47.32 47.18 0.00 0.00 34.06 13.29 41.95 49.72 0.00 0.00 41.66 35.69
SumAggGIN (Sheng et al., 2020) 13.26M 122.69s 78.19 77.82 52.27 54.11 2.17 2.31 35.79 36.43 54.15 55.07 4.05 2.12 48.31 47.22
DisGCN (Sun et al., 2021) 13.17M 108.71s 70.84 76.67 42.71 46.13 1.17 1.55 32.08 16.97 50.03 50.17 2.35 1.99 38.25 39.97
MM-DFN (Hu et al., 2022a) 5.33M 45.63s 78.17 77.76 52.15 50.69 0.00 0.00 25.77 22.93 56.19 54.78 0.00 0.00 48.31 47.82
M2FNet (Chudasama et al., 2022) 8.15M 66.54s 72.88 67.98 72.76 58.66 5.57 3.45 50.09 47.03 68.49 65.50 17.69 25.24 57.33 55.25
EmoCaps (Li et al., 2022b) 12.31M 101.77s 75.24 77.12 63.57 63.19 3.45 3.03 43.78 42.52 58.34 57.05 7.01 7.69 58.79 57.54
LR-GCN (Ren et al., 2021) 14.97M 147.38s 81.51 80.83 55.42 57.11 0.00 0.00 36.36 36.96 62.21 65.84 7.32 11.07 52.63 54.74
@@
To comprehensively evaluate the practicality of different MCER models, we compare their parameter sizes and inference time in addition to their recognition performance. Tables 8 and 9 report the number of parameters and average running time per inference on the IEMOCAP and MELD datasets, respectively. The results show that lighter models such as TextCNN and bc-LSTM possess significantly fewer parameters (0.47M and 1.28M, respectively) and exhibit faster inference times (0.96s and 2.16s). However, their performance across most emotion categories tends to be lower compared to more complex architectures. In contrast, recent transformer-based or graph-enhanced models, such as DialogueGCN, MM-DFN, M2ETNet, and especially LR-GCN, require substantially more parameters (up to 15.77M) and longer inference times (up to 21.06s for IEMOCAP and 147.38s for MELD), but consistently achieve better recognition performance. This observation highlights a common trade-off between model complexity and computational efficiency. While heavier models are better suited for applications where accuracy is the priority, lightweight models may be more favorable for real-time or resource-constrained deployment scenarios. In particular, LR-GCN achieves the best overall performance on both datasets, ranking highest in multiple emotion categories, albeit with the largest parameter count and slowest inference time. This suggests that while model complexity improves expressive power, there is a pressing need to explore model compression, pruning, and quantization to improve the feasibility of deploying such models in real-world applications.
@@ 7.1. Social Media Analysis
Multi-modal conversational emotion recognition has many broad applications in social media analysis (Zhang et al., 2021). The most typical application is product improvement and innovation, that is, by analyzing user comments and feedback on social media, companies can understand users’ preferences and dissatisfaction with products. This helps companies tweak product designs, improve functionality, and develop products that better meet user needs. Therefore, businesses can employ emotion analysis techniques to improve their products. In addition, emotion analysis can also help advertisers understand users’ emotional attitudes towards advertisements, thereby optimizing advertisement content and strategies, and improving advertisement effectiveness. For example, FaceReader (Li and Zhang, 2024) can measure people’s emotional responses to different advertising creatives, providing valuable insights into the effectiveness of emotional appeals, humor, or shock value. With an 89% recognition rate for static images and an 80% recognition rate for animated expressions, FaceReader provides a reliable method for assessing the emotional impact of advertising imagery.